NDOS Commissioning Meeting 2011/03/28¶
Attendance: Jon, Andrew, Gary, Sue, Mark Mathis, Alex Sousa, Mat, Denis, Rick K., Alena, Holger, Minerba, Jarek, Gavin, Leon, Art
- CRL -> ECL move
- NDOS Status
- DAQ/Electronics Update & Plans
- DCS Status & Plans
CRL -> ECL move:¶
- Looks like a smooth transition...
- Both Denis and Alex have some suggestions
- Jon will disable everything except for General
- Leon: users manual!?
- Mat: scintillator filling completed for large portion of 3rd diblock and muon catcher. This includes everything that could be reached with filling nozzle. Karen is modifying the filling head to be able to reach some of the modules. She'd like to try to fill some empty modules in the first 2 di-blocks today.
- Jon will add some How-To information to the wiki re: powering off individual systems.
- There will be a test of the muon-catcher's water system today.
- FEB installation will begin in a couple of days. Will start with top of diblock 3.
- USB connections at NDOS: partially installed, the cables are run and the hubs are mounted in rack 2. Ready to try connecting to DCMs.
- tried grounding heatsinks to the clip, and it didn't seem to have any effect on the noise.
- Gerard noticed that noise is correlated to how tight the box cover is on. This will be checked during the cap-board installation.
- Progress on DCM WOFO? Sue has not been able to reproduce the problem on the teststand, but that was before they realized that a different version of TDU firmware was being used on the teststand than on the detector. Hasn't had a chance to go back to try again.
- Sue will be updating Ganglia interface shortly after this meeting. Updates will include ability to select time to display, ability to display many reports for a cluster.
- Leon: DCM firmware that didn't go through on Friday? Will happen later today (after lunch)
- Jarek: noticed while on shift that data taking is very smooth when no on is in the NDOS building.